# SurGE: A Benchmark and Evaluation Framework for Scientific Survey Generation

### v2

@@ Abstract
The rapid growth of academic literature makes the manual creation of scientific surveys increasingly infeasible. While large language models show promise for automating this process, progress in this area is hindered by the absence of standardized benchmarks and evaluation protocols. To bridge this critical gap, we introduce SurGE (Survey Generation Evaluation), a new benchmark for scientific survey generation in computer science. SurGE consists of (1) a collection of test instances, each including a topic description, an expert-written survey, and its full set of cited references, and (2) a large-scale academic corpus of over one million papers. In addition, we propose an automated evaluation framework that measures the quality of generated surveys across four dimensions: comprehensiveness, citation accuracy, structural organization, and content quality. Our evaluation of diverse LLM-based methods demonstrates a significant performance gap, revealing that even advanced agentic frameworks struggle with the complexities of survey generation and highlighting the need for future research in this area. [^1] We have open-sourced all the code, data, and models at: https://github.com/oneal2000/SurGE.
@@ 1 Introduction
The volume of scientific literature has been expanding at an unprecedented rate in recent years. For instance, academic archives like arXiv now receive over a thousand new computer science papers daily, more than doubled between 2019 and 2024 (Liang et al., 2025). This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as the manual collection and synthesis of large volumes of relevant papers is both labor-intensive and time-consuming. Faced with this challenge, there is a growing need for automated systems that can generate survey papers effectively. While recent advancements in LLM-based agents (Yao et al., 2023; Wang et al., 2024c) offer significant promise for automating this task, their full potential is severely hindered by the lack of reliable, scalable, and standardized evaluation benchmarks. Currently, the evaluation of automatically generated surveys largely depends on human-expert reviews, which limit the reproducibility and objectivity of assessments (Tian et al., 2024a). Consequently, progress in this field is difficult to quantify, and the relative merits of different methods remain challenging to compare without standardized, multi-faceted evaluation benchmarks.
@@
To address these challenges and fill this research gap, we introduce SurGE (Survey Generation Evaluation), a novel benchmark that establishes a reproducible standard for the survey generation task, which we formalize as a two-stage process: (1) retrieving relevant papers for a given topic from a large corpus, and (2) synthesizing them into a well-structured survey. To support standardized evaluation of both stages, SurGE provides a comprehensive dataset consisting of two key components. The first is a large-scale academic corpus containing over one million computer science papers to be used as the search and retrieval pool. The second is a collection of test instances, where each instance comprises a research topic (e.g., ”Machine Learning for Information Retrieval”) and its corresponding ground-truth survey. Each ground-truth survey is a high-impact, peer-reviewed paper, chosen based on strict citation criteria and subsequently validated for its quality by expert annotators (see Section 3.1), and is provided along with its full list of cited references.
Beyond dataset construction, another key contribution of this work is a fully automated evaluation framework that enables reproducible and multi-faceted assessment of generated surveys. Grounded in established principles for high-quality survey writing (Webster & Watson, 2002; Boote & Beile, 2005; Keele et al., 2007; Pautasso, 2013), our proposed evaluation framework quantifies a survey’s quality across four key dimensions: comprehensiveness, citation accuracy, structural organization, and content quality. This automated approach eliminates the reliance on subjective and time-consuming human evaluation, thus establishing a scalable and standardized methodology to guide and measure progress in survey generation.
To validate the utility of SurGE, we evaluate a range of LLM-based baseline systems on the benchmark. These include standard retrieval-augmented generation approaches that first retrieve relevant documents and then generate a survey, as well as more advanced pipelines that incorporate explicit planning (e.g., outline generation) and iterative refinement of the draft. The results reveal that even the state-of-the-art systems struggle with the survey generation task. For example, models often miss important papers, produce fragmented or imbalanced coverage of subtopics, and sometimes generate hallucinations with irrelevant citations. These findings indicate a significant gap between machine-generated surveys and expert-written ones, underscoring the need for further research in this area. We believe SurGE will spur the development of more effective techniques at the intersection of information retrieval and generative modeling to tackle this challenging task.
@@
    We introduce SurGE, a comprehensive benchmark for scientific survey generation, featuring expert-written ground-truth surveys and a large-scale academic paper corpus.
@@
    We provide extensive baseline results and analyses, offering reference points and highlighting key challenges in this emerging task.
@@ 2 Task Definition
We formalize scientific survey generation as a two-stage task. Given a topic description $t$ and a large academic corpus $D=\{d_{1},d_{2},\ldots,d_{n}\}$, the goal is to automatically generate a survey article $S$ that provides a structured and comprehensive overview of the topic. The process consists of:
  • Document Collection: A retrieval module or complex agentic retrieval system collects a topic-relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
  • Survey Generation: A generative model composes a well-structured survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a reference list.
To enable standardized benchmarking of the survey generation task, SurGE offers a comprehensive framework that formalizes the process. The benchmark includes 205 expert-selected topics, each paired with a topic description $t$ and a ground-truth survey for evaluation. SurGE also provides a large-scale academic corpus $D$ containing over one million papers, which supports the document retrieval stage. The generated surveys are then evaluated through an automated system, which quantitatively assesses their quality across four key dimensions. Details of the dataset construction process are outlined in Section 3, while the evaluation framework is explained in Section 4.
@@ 3 Dataset Construction
3.1 Ground-Truth Survey Collection and Expert Annotation
To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature. Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews. To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann & Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
Following the initial filtering process, we further refine the SurGE dataset through expert annotation. This process aims to assess not only the citation-based impact of each candidate survey but also its quality from the perspective of experienced researchers. To this end, we recruited a team of four computer science Ph.D. students as annotators. Note that annotating a paper did not require a close reading of the entire document. Annotators could typically complete the task within 8–9 minutes per paper on average. Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) citation quality, which examines the relevance, diversity, and traceability of cited works.
Each annotator labeled the document as either “usable” or “not usable.” A survey paper was included in the final dataset only if both annotators independently marked it as usable. In cases of disagreement, the paper was discarded to maintain a conservative quality threshold. To ensure fair and motivated participation, annotators were compensated based on working hours at an average rate of 60 CNY per hour, exceeding the local minimum wage in Beijing. Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances. The resulting score of 0.792 indicates substantial agreement and reinforces the reliability of the quality control process. After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers. The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix F. For each of the 205 selected ground-truth surveys, we provide its complete reference list, which serves as the gold standard for evaluating the comprehensiveness of a generated survey. The detailed methodology for extracting and processing these references is described in Appendix A.
@@ 3.2 Academic Corpus Construction
The corpus was constructed through a two-stage process. The initial stage involved seeding the corpus with the references from our 205 ground-truth surveys. We systematically retrieved the arXiv metadata for every cited paper that was publicly accessible. This process revealed that approximately 30% of the references were unavailable, primarily due to publication in closed-access journals or other restricted venues. In the second stage, we expanded the corpus to enhance its comprehensiveness. We queried the official arXiv search API, using keywords and titles from the ground-truth surveys to identify and collect metadata for other topically related papers. This methodology resulted in a final retrieval corpus of 1,086,992 unique papers. For each paper, the corpus provides rich metadata, including its title, authors, abstract, subject categories, publication date, and a direct link to its arXiv page for transparency and verification. To ensure high data quality, all collected metadata underwent a rigorous pre-processing pipeline involving text normalization, formatting removal, and deduplication.
@@ 3.3 Statistics and Analysis
The resulting SurGE benchmark comprises 205 ground-truth survey papers and a retrieval corpus of 1,086,992 documents. Table 1(a) presents the key statistics of our curated dataset. To quantitatively analyze the organizational complexity of the surveys, we model the hierarchical section headings (e.g., Section, Subsection) of each paper as a tree structure. Our analysis reveals that these surveys are structurally deep, with an average tree depth of 3.07 and a mean of 42.7 nodes (i.e., distinct sections) per document. This structural complexity presents a significant challenge for hierarchical text generation. Furthermore, the surveys are densely referenced, citing an average of 65.8 papers, which underscores the demand for high-recall information collection. Table 1(b) details the pre-processed fields for each survey instance, which include not only standard metadata but also the parsed structural tree and the ground-truth list of cited documents, enabling a fine-grained, multi-faceted evaluation of system-generated surveys.
@@ 3.4 Ethical Considerations and Licensing
Our corpus is constructed exclusively from arXiv-provided descriptive metadata (titles, authors, abstracts, identifiers, categories, and license URIs) harvested via the official API. We do not host or redistribute arXiv PDFs or source files. This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication (arx, 2025b). This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata (arx, 2025a).
We have prioritized transparency, reproducibility, and ethical considerations throughout dataset construction. To support open science, we have publicly released the SurGE dataset, accompanying metadata, and all associated processing scripts on our official GitHub repository [^2] https://github.com/oneal2000/SurGE. The dataset and codebase are distributed under the MIT license, granting researchers and developers unrestricted access and modification rights. Regular updates will ensure continued relevance and alignment with evolving research trends and standards.
Figure 1: Overview of the SurGE Benchmark. (a) Summary statistics of the curated survey dataset and its associated retrieval corpus. (b) Metadata of the pre-processed survey dataset used in SurGE.
(a) Basic Statistics of the SurGE Dataset and Corpus

### v4

@@ Abstract
The exponential growth of scientific literature has created a pressing need for automated survey generation. Although recent LLM-based agents have shown promise in automating this task, current progress is hindered by the lack of a standardized, scalable evaluation protocol. Existing evaluation methods typically rely on either human evaluation or custom metrics designed to validate specific pipelines, which restricts scalability and hinders fair comparison. To address this, we introduce SurGE, a benchmark and evaluation framework tailored for scientific survey generation. SurGE provides a large-scale retrieval corpus of over one million papers and expert-validated ground-truth surveys. Furthermore, we propose a robust multi-dimensional evaluation protocol that integrates both objective metrics and LLM-based judgments, and empirically verify its high alignment with human experts. Our experiments reveal that while agentic pipelines outperform RAG baselines in fluency and structural quality, they still struggle with citation accuracy, highlighting key directions for future research [^1] We have open-sourced all the code and data at: https://github.com/oneal2000/SurGE.
SurGE: A Benchmark and Evaluation Framework for
Scientific Survey Generation
@@ 1 Introduction
The volume of scientific literature has been expanding at an unprecedented rate in recent years. Academic archives like arXiv now receive over a thousand new computer science papers daily, a figure that has more than doubled between 2019 and 2024 (Liang et al., 2025). This rapid growth in publications has made the manual creation of comprehensive survey papers increasingly impractical, as collecting and synthesizing such vast volumes of information is both labor-intensive and time-consuming. Consequently, there is a critical need for automated systems capable of generating high-quality academic surveys.
Recent LLM-based agents have shown promise in automating this task through hierarchical decomposition or memory-driven mechanisms (Wang et al., 2024; Yan et al., 2025; Lai et al., 2024). However, while the feasibility of generation has been established, the field currently lacks a standardized evaluation protocol. Existing evaluation methods are often unscalable or strictly tied to specific generation pipelines. For example, StepSurvey relies heavily on human evaluation (Tian et al., 2024a), which severely restricts scalability and reproducibility. Conversely, framework-oriented works like AutoSurvey and SurveyForge propose custom metrics tailored to validate their own pipelines. This "self-validation" paradigm inevitably introduces bias and hinders fair, cross-system comparison. As the community moves beyond "making it work" to "making it work well," there is a critical need for a method-agnostic, reproducible, and diagnostic benchmark.
@@
To address this gap, we introduce SurGE (Survey Generation Evaluation), a comprehensive benchmark and evaluation framework designed to serve as a universal yardstick for the survey generation task. We formalize this task as a two-stage process: (1) retrieving relevant papers from a massive corpus, and (2) synthesizing them into a structured survey. To support this, SurGE provides a rigorously constructed dataset consisting of a large-scale academic corpus (over one million papers) and a collection of high-quality test instances. The ground-truth surveys are sourced from high-impact, peer-reviewed publications, and are also strictly validated by expert annotators to ensure data reliability and ethical compliance (see Section 3.1).
Beyond dataset construction, a key contribution of SurGE is its robust, multi-dimensional evaluation framework specifically tailored for survey generation. Grounded in established principles for high-quality survey writing (Webster and Watson, 2002), we propose a comprehensive evaluation framework that integrates objective, quantifiable indicators (e.g., citation coverage) with LLM-as-a-Judge mechanisms (e.g., assessing logical coherence and content quality). Crucially, to ensure the reliability of our proposed evaluation framework, we conduct a rigorous meta-evaluation: we collect fine-grained expert annotations and verify a high correlation between human judgment and our LLM-based metrics. This alignment confirms that SurGE serves as a trustworthy and scalable proxy for the evaluation of AI-generated surveys.
To empirically assess the utility of SurGE, we benchmark diverse LLM-based systems. Leveraging the decoupled retrieval-generation design of our framework, we conduct a granular analysis to disentangle whether performance bottlenecks arise from information acquisition or content synthesis. Our experiments demonstrate that specialized agentic pipelines consistently outperform standard RAG baselines. However, while these agents exhibit superior linguistic fluency, they struggle with information utilization, failing to fully leverage the collected evidence during the generation phase.
@@
    We introduce SurGE, a method-agnostic benchmark for survey generation featuring expert-validated ground-truth surveys and a large-scale retrieval corpus.
@@
    We conduct extensive experiments on representative state-of-the-art systems, pointing out specific directions for future optimization.
@@ 2 Task Definition
We formalize survey generation as a two-stage task. Given a topic description $t$ and a large academic corpus $D=\{d_{1},d_{2},\ldots,d_{n}\}$, the goal is to automatically generate a survey article $S$ that provides a structured and comprehensive overview of the topic. The process consists of:
  • Information Collection: A retrieval system collects a relevant paper set $\mathcal{R}_{t}\subseteq D$ containing papers relevant to $t$.
  • Survey Generation: A generative system composes a survey $S$ based on the topic $t$ and the retrieved document set $\mathcal{R}_{t}$, including proper citations and a proper reference list.
@@ 3 Dataset Construction
3.1 Ground-Truth Survey Collection
To construct the SurGE benchmark, we began by collecting a diverse set of high-quality reference surveys (ground-truth surveys) from recent computer science literature. Candidate texts were drawn from the arXiv repository, focusing on publications between 2020 and 2024 that self-identified as survey articles or systematic reviews. To ensure the academic significance and reliability of each instance, we applied the following selection criteria: (i) the document must explicitly declare itself as a survey or review; (ii) it must have achieved a minimum citation count of 20, indicating scholarly impact (Bornmann and Daniel, 2008); and (iii) the publication date must be between 2020 and 2024.
3.2 Expert Annotation
Building upon the initial automated collection, we implemented a secondary manual filtering phase to ensure the quality of the dataset. To this end, we recruited a team of four Ph.D. students in computer science, led by a faculty member, to serve as annotators. Each candidate document was evaluated by two independent annotators along four key dimensions: (i) citation impact, reflecting the scholarly influence of the paper; (ii) content coverage, indicating how comprehensively the survey summarizes the literature within its scope; (iii) structural coherence, assessing the logical organization and clarity of the document’s sections; and (iv) citation quality, which examines the relevance, diversity, and traceability of cited works.
Annotation averaged 10 minutes per paper. Each annotator labeled the document as either “usable” or “not usable”. Papers were retained only if independently labeled as “usable” by both annotators, discarding any disagreements to maintain a conservative quality threshold. To ensure motivation, annotators were compensated 60 CNY/hour (much exceeding the local minimum wage). Inter-annotator agreement was quantified using Cohen’s Kappa, applied to 250 annotated instances. The resulting score of 0.792 indicates substantial agreement, reinforcing the reliability of the quality control process. After this filtering stage, we finalized the dataset with 205 rigorously verified survey papers. The final set of 205 ground-truth surveys included in our benchmark is detailed in Appendix H.
@@ 3.3 Academic Corpus Construction
The corpus was constructed through the following process. First, we retrieved the arXiv metadata for all publicly accessible papers cited in the ground truth survey. This process revealed that around 30% of the references were unavailable, primarily due to publication in closed-access journals or other restricted venues. Then we expanded the corpus to enhance its comprehensiveness. We queried the official arXiv search API, using keywords and titles from the ground-truth surveys to identify and collect metadata for other topically related papers. To mitigate topic bias, we augment the corpus with random sample of arXiv computer science papers, drawn independently of the benchmark topics. The entire process resulted in a final retrieval corpus of 1,086,992 unique papers. For each paper, the corpus provides rich metadata, including its title, authors, abstract, subject categories, publication date, and a direct link to its arXiv page for transparency and verification. To ensure high data quality, all collected metadata underwent a rigorous pre-processing pipeline involving text normalization, formatting removal, and deduplication.
@@ 3.4 Statistics and Analysis
The resulting SurGE benchmark comprises 205 ground-truth survey papers and a retrieval corpus of 1,086,992 documents. Table 1 presents the key statistics of the dataset. We model hierarchical headings as trees to quantify the complexity of the ground truth surveys. The surveys exhibit deep structures (avg. depth 3.07, 42.7 nodes), posing significant challenges for hierarchical text generation. Furthermore, the surveys are densely referenced, citing an average of 65.8 papers, which underscores the demand for high-recall information collection. Detailed descriptions of the metadata fields for both the survey instances and the retrieval corpus are provided in Appendix B.
@@ 3.5 Ethical Considerations and Licensing
Our corpus is constructed exclusively from arXiv-provided descriptive metadata (titles, authors, abstracts, identifiers, categories, and license URIs) harvested via the official API. We do not host or redistribute arXiv PDFs or source files. This design complies with arXiv’s API Terms of Use, which place descriptive metadata under a CC0 public-domain dedication (arx, 2025b). This design is also consistent with the arXiv Submittal Agreement’s CC0 designation for metadata (arx, 2025a). A more detailed discussion of ethical considerations is provided in Section §10. To support open science, we have publicly released the SurGE dataset, accompanying metadata, and all associated processing scripts on our official GitHub repository. The dataset and codebase are distributed under the MIT license, granting researchers and developers unrestricted access and modification rights.
Table 1: Key statistics of the SurGE benchmark dataset and its associated retrieval corpus.